A Bibliography of Publications in IEEE Parallel and Distributed Technology: Systems and Applications

Nelson H. F. Beebe Center for Scientific Computing

University of Utah

Department of Mathematics, 322 INSCC

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

FAX: +1 801 585 1640, +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@ieee.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

01 March 2018 Version 1.14

Title word cross-reference

1 [Mes93].

2 [Bok96].

5 [Hed93].

90D [PH96]. ’94 [Qui95b]. ’96 [Ano96t].

Abstraction [AFK^{+}93]. Achieving [MFS94]. Across [HM94]. action [Fie94]. Adaptive [BCSL95, KOR95, NX93, KKM93]. Advanced [CZM94]. Advertiser [Ano95b, Ano95c, Ano95d, Ano96c, Ano96d, Ano96e]. Advertiser/Product [Ano95b, Ano95c, Ano95d, Ano96c, Ano96d, Ano96e]. Affects [LHWS96]. age [Hed96b, Ano96v]. ahead [Qui94c]. algebra [Don93]. Algorithm [CH94]. Algorithms [ZYC96, GGK93]. Analysis [SMB95]. Analyzing [BBV95, LW94]. Ancient [Tay96a]. Apertos [Tok93]. Application [FYC95]. Applications [BCDP95, BCSL95, CZM94, CFS95, Rov93, SSLP93, WB94]. Applied [Lev94]. Approach [Du93, HRE^{+}95, RHWG95]. Approaches [GTM96]. apps [Bel94]. architecture [Pfi96]. Architectures [Alm96, BCSL95, CO94, GY93, GGK93]. Argonne [DEH^{+}95]. Article [Ano96f, Ano96g, Ano96h, Ano96i]. Assess [WB96]. Assessing [MABG96]. ATM [HM94]. Automated [SMB95]. Automatic [BEH^{+}94]. Automating [BBH^{+}94].

back [Ano96-33, Ano96-34, Qui94c]. Balancing [LHWS96, MW95]. Based [GGJ95, HRE^{+}95, Hin95]. Basic [TWRP96]. Batrun [TKDA96]. be [Bel94]. benchmark [BBDS93]. Benchmarking [ZFV96]. Beyond [DC93, PC94]. Book [Ano95e, Ano95f, Ano95g, Ano95h, Ano96j, Ano96k, Ano96l]. Books [Ano95n, Ano96w, Ano96x]. Bounded [Gro93]. breakthroughs [WB94]. Broadcast [Kah96]. Butterfly [ZCC94]. Buy [Cla95].

C [Bur93]. Cache [CLY96, HKSL96]. Calendar [Ano95i, Ano95j, Ano96m, Ano96n, Ano96o, Ano96p, Tay95b, Tay95c, Tay95d]. Calendar/Call [Ano95i, Ano95j, Ano96m, Ano96n, Ano96o, Ano96p, Tay95b, Tay95c, Tay95d]. call [Fie94]. Cards [Ano95u, Ano95v, Ano95w, Ano95x, Ano96-32, Ano96-33, Ano96-34]. Carnot [TCMW93]. Case [SZ96, Zom96]. Cellular [Lin96]. Challenge [BEH^{+}94]. Challenges [Agh95b]. changing [Pan93]. Characterization [CMM^{+}95]. Checkpointer [PL94]. Chief [Agh95a, Agh95b, Agh96b, Agh96c, Agh96a, Ano96u, Qui94c, Qui94a, Qui94b]. Clusters [Pfi96]. CM [Hed93]. CM-5 [Hed93]. Code [MWO95, Wei93]. Coherence [CLY96, MC94]. Collaborative [DEH^{+}95]. Combating [Hed96a]. commercial [Pfi96]. Communication [Hin95, HM94, MABG96, XH96]. Communications [CDH^{+}95]. compare [Cro94]. Compile [PH96]. Compile-Time [PH96]. Compiler [CLY96, Spo94]. Compiler-Directed [CLY96]. Compiling [GY93, PAM94]. Complete [Bok96, RHWG95]. Components [BCDP95]. computation [ESM^{+}94].

Computational [KOR95]. Computer [RHWG95, Hed96a, WB94, Ano95k, Ano95l, Ano95m, Ano96q, Ano96r, Ano96s]. computers [Don93, Pfi96, Wal96]. Computing [BEH^{+}94, BHD^{+}95, CFS95, HM94, RS94, TKDA96, AFK^{+}93, Cla95, Hil93, Kah96, Wal96, Wil93, Per96]. Concepts [PTM96]. Concurrent [TWRP96, AFK^{+}93, Mes93, YMYT93]. Conferences [Ano96t, Qui95b, RCC^{+}95, SSBK94]. Connection [Hin95]. Connection-Based [Hin95]. Consistent [PL94]. consortium [Mes93]. constructs [LW94]. Control [BBH^{+}94, SD94]. cover [Ano96-33, Ano96-34]. Cowichan [WB96]. Cray [MWO95]. Cruising [SSBK94]. CS [Bok96]. CS-2 [Bok96].

Data

[ACG^{+}94, CZM94, MC94, PHD^{+}95, Ree95]. Data-Parallel [ACG^{+}94, CZM94, PHD^{+}95]. Dataflow [HKSL96]. DAVE [FYC95]. DBS3 [ZFV96]. Debugger [SAB^{+}94, MB93]. Debugging [BHD^{+}95]. Decentralized [LML^{+}93]. Dedos [HLV^{+}94]. Defining [LW94]. Demand [JC95, TCKK95]. Design [BCDP95]. Designing [DEH^{+}95, JC95, ZYC96]. Detection [BEH^{+}94]. developing [SSLP93]. Development [BBH^{+}94, FYC95, MT96, LML^{+}93]. dinosaur [Tay95a]. Directed [CLY96]. directions [Ano96u]. disappearance [Tay95a]. Dispatchers [KKD94]. Distributed [BCDP95, BBH^{+}94, CF93, CH94, ERLS93, ESM^{+}94, FYC95, MFS94, PSY95, PTM96, RS94, SD94, Tay96b, Cla95, HLV^{+}94, KKM93, PAM94, SSLP93, Tok93]. distributed-memory [PAM94]. distributing [Bur93]. Distributions [PHD^{+}95]. DSM [GTM96]. Dynamic [BCSL95, GSN93, LHWS96]. economic [Wal96]. Editor [Agh95a, Agh95b, Agh96b, Agh96c, Agh96a, Alm96, Ano96u, HCW94, Hat94, Pan93, Qui94c, Qui94a, Qui94b]. Editor-in [Qui94b]. Editor-in-Chief [Agh95a, Agh95b, Agh96b, Agh96c, Agh96a, Ano96u, Qui94c, Qui94a]. Editors [CFS95, PSY95, RS94]. elusive [Agh96b]. endangered [Ano96v, Hed96b]. Enhanced [CDH^{+}95]. Enterprise [SSLP93]. Environment [WFL96, HLV^{+}94, Vee94]. Environments [ACG^{+}94, KOR95, PSM93]. Evaluating [ZYC96]. Evaluation [PSY95]. Exchange [Bok96]. Exclusion [ZYC96]. Execution [PC94, RP93]. Exhibition [Qui95b]. Expanding [PC94]. Experiments [MW95]. Exploiting [MC94, YLB93]. ExSpect [BBV95]. Extending [CZM94]. extensible [MB93, TCMW93].

face [Agh96a, Pan93]. false [Gus94]. File [KN95]. File-System [KN95]. Flexible [SD94]. Floor [Qui95b]. Focus [ST96, Tay95a, Tay96a, Tay96b, Wal96]. Fortran [GOS94, Hat94, Lov93, MWO95]. Fortran/PVM [MWO95]. Framework [GOS94]. Free [Wal96]. Free-market [Wal96]. frontier [WB94]. functional [PAM94]. FunRun [SSBK94]. Future [DEH^{+}95, Hil93]. Futures [DEH^{+}95].

Generation [MW95]. Getting [RCC^{+}95]. Global [Kah96, Gro93, Wal96]. glossary [Wil93]. goal [Agh96b]. goals [Gus94]. Grand [BEH^{+}94]. Graph [TWRP96]. Graphical [BCDP95]. Graphs [KOR95]. Group [Sch94]. Growing [Agh95b]. Guest [Alm96, CFS95, HCW94, Hat94, Pan93, PSY95, RS94]. Guide [Per96].

Hardware [GTM96]. Hardware-Oriented [GTM96]. Heterogeneous [MFS94]. High [BEH^{+}94, CDH^{+}95, CFS95, Fos94, JC95, Lov93, Don93, GOS94, Hat94]. High-Performance [BEH^{+}94, CFS95, Fos94, JC95, Don93]. High-Speed [CDH^{+}95]. HPC [Fie94]. HPF [CZM94, PH96, Spo94, Van94, Vee94, WSL94]. HPF/Fortran [PH96]. Hybrid [HRE^{+}95].

I/O [FCBH95, RCC^{+}95]. IBM

[Hed96a, MABG96, WFL96, XH96]. ickp [PL94]. Idle [TKDA96]. immune [Hed96a]. Impact [Hat94]. Implementing [JC95, YMYT93]. Improve [MC94]. incremental [NX93]. Index [Ano95b, Ano95c, Ano95a, Ano96c, Ano96d, Ano96e, Ano96b, Ano95d]. India [Kah96]. Industry [Ano96v, Hed96a, Hed96c, WB94]. Info [Ano95k, Ano95l, Ano95m, Ano96q, Ano96r, Ano96s]. information [Hed96b, Ano96v]. infrastructure [Wal96]. Inherently [KKD94]. instrumentation [Rov93]. integration [Wal96]. intelligence [Agh96b]. Intensional [Du93]. Interlockings [BBV95]. Introduction [Alm96, CFS95, HCW94, Hat94, PSY95, RS94, Pan93]. invisible [Pfi96]. IPPS [Ano96t]. Irregular [PHD^{+}95, TWRP96]. Isoefficiency [GGK93]. Issues [HM94].

Keeping [ST96]. KSR1 [ZCC94].

labor [Agh96c]. Laboratory [DEH^{+}95]. LAN [CDH^{+}95]. Languages [Fos94, PHD^{+}95, YMYT93]. laptops [Wal94]. Large [TKDA96]. Large-Scale [TKDA96]. leap [Cla96]. Learned [MWO95]. LeMans [SSBK94]. Lessons [MWO95]. libraries [Don93]. Linear [Don93]. List [KKD94]. Load [LHWS96, MW95, KKM93]. lock [ZCC94]. logging [NX93]. Looking [Qui94c].

Machine [Per96, SZ96, Wei93]. machines [Ano96t, WSL94]. magazine [Agh96a].

Magic [TCKK95]. main [Tay95a]. Management [Lin96]. Managing [CLFZ95]. Manipulators [SD94]. Mapping [GY93]. market [Wal96]. massive [Ano96t]. Massively [BCSL95, FCBH95, SAB^{+}94, Ano96t]. measure [Cro94]. measuring [GGK93]. mechanisms [AFK^{+}93]. Medea [CMM^{+}95]. Media [JC95]. Media-on-Demand [JC95]. Membership [Ano95k, Ano95l, Ano95m, Ano96q, Ano96r, Ano96s]. Memera [Hed93]. Memories [HKSL96]. Memory [PTM96, SZ96, ZYC96, Hed93, PAM94]. Merging [CH94]. Message [CO94, SMB95, SAB^{+}94, NX93]. Message-Passing [CO94, SMB95, SAB^{+}94]. Metrics [ST96]. Microsoft [Cla95]. millennium [Hed96c]. Minimizing [MT96]. minimum [Gro93]. mining [Ree95]. Mobility [Lin96]. Model [FYC95, PSM93, SSLP93]. Modeling [GGJ95, XH96]. Models [PC94, ESM^{+}94]. modern [Tay96a]. modularity [AFK^{+}93]. Monitoring [HRE^{+}95]. mountains [Agh96c]. MPI [XH96]. MPL [XH96]. MPPs [Bel94, Cow94]. Multiapplication [IM94]. multicomputer [Rov93]. Multicomputers [PL94, YMYT93]. Multimedia [CFS95, FYC95, Tay96b]. Multiphase [Bok96]. Multiple [SSHT93]. Multiprocessor [KN95]. Multiprocessors [ZYC96]. Mutual [ZYC96].

naming [RP93]. NAS [BBDS93]. nets [CF93]. Network [LHWS96, Hed96c, DC93]. Networked [Per96]. Networks [CLFZ95, HM94, Lin96]. News [Ano95s, Ano95t, Ano96-29, Ano96-30, Ano96-31]. next [WB94]. Nonuniform [KOR95]. number [Ano96t].

O [FCBH95, RCC^{+}95]. Object [GGJ95, Bur93, GSN93, Tok93, YMYT93].

Object-Based [GGJ95]. object-oriented [GSN93, Tok93, YMYT93]. Oklahoma [SSHT93]. operating [Tok93]. optimal [LML^{+}93]. Optimizer [ZFV96]. Optimizing [Wei93]. Orca [WB96]. Oriented [GTM96, GSN93, Tok93, YMYT93]. other [Gus94]. Our [Agh96a]. Overhead [MT96, XH96].

Pacific [Van94]. panel [Ano96t]. Panorama [MB93]. Papers [Ano95i, Ano95j, Ano96m, Ano96n, Ano96o, Ano96p, Tay95b, Tay95c, Tay95d]. Paragon [Bok96]. Parallax [LER93]. Parallel [ACG^{+}94, Alm96, BCSL95, BHD^{+}95, CMM^{+}95, CZM94, CO94, Du93, ERLS93, FCBH95, GY93, GGJ95, HCW94, HMR95, HM94, Kah96, Lev94, MC94, Mra94, PSY95, Per96, PHD^{+}95, RCC^{+}95, RS94, SMB95, SAB^{+}94, WB94, ZFV96, Zom96, Ano96t, BBDS93, Cow94, Cro94, GGK93, GSN93, Hil93, IM94, LER93, LML^{+}93, MB93, Wil93, YLB93]. parallel-program [IM94]. Parallelism [BEH^{+}94, Fos94, GOS94, Ree95, PAM94]. Parallelization [MT96]. Partial [MT96]. Partitioning [KOR95]. Passing [CO94, SMB95, SAB^{+}94]. past [Hil93]. 77toHPF [Van94]. Call [Ano95i, Ano95j, Ano96m, Ano96n, Ano96o, Ano96p, Tay95b, Tay95c, Tay95d]. Fortran [PH96]. Product [Ano95b, Ano95c, Ano95d, Ano96c, Ano96d, Ano96e]. PVM [MWO95]. Subscription [Ano95u, Ano95v, Ano95w, Ano95x, Ano96-32, Ano96-33, Ano96-34]. per-process [RP93]. Performance [BEH^{+}94, CFS95, Fos94, GOS94, HRE^{+}95, Hat94, HMR95, JC95, MABG96, PSY95, PH96, PC94, ST96, SZ96, XH96, Cro94, Don93, IM94, Lov93]. Personal [Ano96v, Hed96b, Don93]. perspective [Don93]. Petri [CF93]. PGHPF [Sch94]. Pixel [CH94]. plans [WSL94]. Play [FYC95]. Plug [FYC95]. Plug-and-Play [FYC95]. Point [Mra94]. Point-to-Point [Mra94]. Polygon [CO94]. portable [MB93]. Porting [MWO95]. Portland [Sch94]. power [LML^{+}93]. Practice [ERLS93, HMR95]. Prediction [PH96, SZ96]. Prepare [Vee94]. present [Cro94]. pricing [LML^{+}93]. Prime [RCC^{+}95]. Priority [KKD94]. privacy [Ano96v, Hed96b]. problem [Bel94]. Problems [TWRP96, WB96]. process [RP93]. Processing [Zom96, Cow94, GSN93, Pfi96]. Products [Ano95o, Ano95p, Ano95q, Ano95r, Ano96y, Ano96z, Ano96-27, Ano96-28]. program [IM94, LER93, LW94, LML^{+}93, NX93]. Programming [ACG^{+}94, BHD^{+}95, Du93, Vee94]. Programs [GY93, GGJ95, Mra94, SMB95, SAB^{+}94, MB93]. propagation [ESM^{+}94]. Putting [Fie94]. PVM [CDH^{+}95, ESM^{+}94, Per96]. Pyrros [GY93].

Query [ZFV96].

Railway [BBV95]. Reader [Ano95u, Ano95v, Ano95w, Ano95x, Ano96-32, Ano96-33, Ano96-34]. Ready [RCC^{+}95]. Real [KKD94, Mra94, RS94, SD94, TCKK95, Zom96, HLV^{+}94, Tok93]. Real-Time [KKD94, Mra94, RS94, SD94, TCKK95, HLV^{+}94, Tok93]. Reducing [Mra94]. redundancy [YLB93]. Referees [Ano96a, Qui95a]. remote [RP93]. Rendering [CO94, HCW94, MC94]. replay [NX93]. report [Ano96t]. Requirements [ACG^{+}94]. Research [Lev94]. reservations [SSHT93]. results [BBDS93]. Reusable [BCDP95]. Review [Per96]. Reviews [Ano95e, Ano95f, Ano95g, Ano95h, Ano96j, Ano96k, Ano96l]. ROME [Bur93]. Routing [BCSL95]. RPC [DC93]. run [Wei93]. Runtime [ST96].

Scalability [SMB95, GGK93]. Scalable [GY93, KKM93, SAB^{+}94, SZ96, Rov93]. Scale [TKDA96]. scheduling [LER93]. Science [DEH^{+}95]. Scientific [KN95]. Search [HRE^{+}95]. sell [Ano96t]. Server [TCKK95]. Servers [JC95]. Service [Ano95u, Ano95v, Ano95w, Ano95x, Ano96-32, Ano96-33, Ano96-34]. Service/ Subscription [Ano95u, Ano95v, Ano95w, Ano95x, Ano96-32, Ano96-33, Ano96-34]. services [TCMW93]. Shared [PTM96, SZ96, ZYC96, Hed93]. Shared-Memory [ZYC96]. Shared-Virtual-Memory [SZ96]. sharing [KKM93]. Sierra [Van94]. SimOS [RHWG95]. Simplifying [Hin95]. Simulating [BBV95]. Simulation [GTM96, RHWG95, TCKK95, Zom96, CF93]. slice [Cla95]. snapshots [Gro93]. Snooping [CH94]. Society [Ano95k, Ano95l, Ano95m, Ano96q, Ano96r, Ano96s]. SofTech [Spo94]. Software [BBH^{+}94, Hed93, ZYC96]. song [Tay96b]. sounds [Tay96a]. SP [WFL96]. SP2 [MABG96, Bok96, XH96]. species [Ano96v, Hed96b]. Speed [CDH^{+}95, YLB93]. speeds [Ree95]. Speedup [MFS94]. Spin [ZCC94]. Spin-lock [ZCC94]. Spotlight [Ano96v, Hed96a, Hed96c]. Stable [KKD94]. Standards [HRE^{+}95]. Standards-Based [HRE^{+}95]. Stormy [Tay95a]. Story [Cla96]. Strategic [Ano96u]. Structurally [SD94]. Study [GTM96, SZ96, Zom96]. Subsystem [MABG96]. Subsystems [FCBH95]. Summaries [Ano96f, Ano96g, Ano96h, Ano96i]. Supercomputers [FCBH95]. supercomputing [Mes93, Pan93, PSM93, Qui95b]. Superlinear [MFS94]. superscalar [Wei93]. support [IM94]. Supporting [PHD^{+}95]. suspect [Tay95a]. switch [TCMW93]. synchronization [ZCC94]. Synthetic [MW95]. System [KN95, MFS94, RHWG95, TCKK95, Hed96a, Lev94, Tok93]. Systems [CMM^{+}95, ERLS93, HKSL96, PSY95, PTM96, Bur93, KKM93, PAM94, WFL96, YLB93].

T3D [MWO95]. take [Ano96t]. Task [Fos94, GOS94]. Techniques [CLY96]. technological [Cla96]. Technology [TWRP96, Tay96a]. telecommunications [Hed96c]. Telephony [Lin96]. Teraflops [Gus94, Wal94]. terminology [Wil93]. Theory [ERLS93, HMR95]. there [Bel94]. Thinking [WSL94]. Time [KKD94, Mra94, PH96, PC94, RCC^{+}95, RS94, SD94, TCKK95, Zom96, HLV^{+}94, Tok93]. Tool [CMM^{+}95, IM94, LER93]. Tools [PSY95, WFL96]. Topology [LHWS96]. Toy [Cla96]. Trace [WFL96]. Transfers [Mra94]. transforming [LW94]. transition [Hil93]. Tutorial [Per96].

Unified [WFL96]. Unstructured [KOR95]. update [SSHT93]. Usability [WB96]. Use [PC94]. User [Per96]. Using [PHD^{+}95, SZ96, WB96, ESM^{+}94]. Utilizing [TKDA96].

Variance [Mra94]. VAST [Van94]. VAST-HPF [Van94]. VAST/77toHPF [Van94]. vector [Wei93]. Video [TCKK95]. Video-on-Demand [TCKK95]. View [Qui95b, RP93]. Viewpoint [Cla95, Cla96, Pfi96]. Virtual [CLFZ95, Per96, SZ96, DC93]. viruses [Hed96a]. vision [Cla95]. Visions [Bel94, Cow94, Fie94, Gus94, Wal94, WB94]. Vista [Rov93]. Visual [BHD^{+}95]. Visualization [HMR95, Rov93, PSM93]. Volume [MC94].

wave [ESM^{+}94]. weather [Tay95a]. will [Ano96t]. won’t [Bel94]. work [Fie94]. Workload [CMM^{+}95, KN95, MW95]. Workstations [TKDA96, Hed93]. World [CLFZ95]. Wrestling [Hil93].

xHPF [Lev94].

Year [Mes93].

References

Adve:1994:RDP

[ACG^{+}94] Vikram Adve, Alan Carle, Elana Granston, Seema Hiranandani, Ken Kennedy, Charles Koelbel, Ulrich Kremer, John Mellor-Crummey, Scott Warren, and Chau-Wen Tseng. Requirements for data-parallel programming environments. IEEE parallel and distributed technology: systems and applications, 2(3):48– 58, Fall 1994. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Agha:1993:AMM

[AFK^{+}93] Gul Agha, Svend Frolund, WooYoung Kim, Rajendra Panwar, Anna Patterson, and Daniel Sturman. Abstraction and modularity mechanisms for concurrent computing. IEEE parallel and distributed technology: systems and applications, 1(2):3–14, May 1993. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Agha:1995:EC

[Agh95a] Gul A. Agha. From the Editor-in-Chief. IEEE parallel and distributed technology: systems and applications, 3(3):2–3, Fall 1995. CODEN IPDTEX.

ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1995/pdf/h30002.pdf.

Agha:1995:ECG

[Agh95b] Gul A. Agha. From the Editor-in-Chief: Growing challenges. IEEE parallel and distributed technology: systems and applications, 3(1):3, Spring 1995. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1995/pdf/h10003.pdf.

Agha:1996:ECO

[Agh96a] Gul A. Agha. From the Editor-in-Chief: Our magazine’s new face. IEEE parallel and distributed technology: systems and applications, 4(3):2–3, Fall 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1996/pdf/p3002.pdf.

Agha:1996:ECE

[Agh96b] Gul A. Agha. From the Editorin-Chief: The elusive goal of intelligence. IEEE parallel and distributed technology: systems and applications, 4(2):2, Summer 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1996/pdf/p2002.pdf.

Agha:1996:ECM

[Agh96c] Gul A. Agha. From the Editor-in-Chief: The mountains are in labor. IEEE parallel and distributed technology: systems and applications, 4(1):2, Spring 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1996/pdf/p1002.pdf.

Almasi:1996:GEI

[Alm96]  George Almasi.  Guest Edi-

tor’s introduction: Parallel architectures. IEEE parallel and distributed technology: systems and applications, 4(4):10, Winter 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1996/pdf/p4010.pdf.

Anonymous:1995:I

[Ano95a] Anonymous. 1995 index. IEEE parallel and distributed technology: systems and applications, 3(4):81–84, Winter 1995. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1995/pdf/h40081.pdf.

Anonymous:1995:APIa

[Ano95b] Anonymous. Advertiser/product index. IEEE parallel and distributed technology: systems and applications, 3(1):96, Spring 1995. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Anonymous:1995:APIc

[Ano95c] Anonymous. Advertiser/product index. IEEE parallel and distributed technology: systems and applications, 3(4):80, Winter 1995. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Anonymous:1995:APId

[Ano95d] Anonymous. Advertiser/product index. IEEE parallel and distributed technology: systems and applications, 3(4):104, Winter 1995. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Anonymous:1995:BRa

[Ano95e] Anonymous. Book reviews.

IEEE parallel and distributed technology: systems and applications, 3(1):89–??, Spring 1995. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1995/pdf/h10089.pdf.

Anonymous:1995:BRb

[Ano95f] Anonymous. Book reviews.

IEEE parallel and distributed technology: systems and applications, 3(2):79–??, Summer 1995. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1995/pdf/h20079.pdf.

Anonymous:1995:BRc

[Ano95g] Anonymous. Book reviews.

IEEE parallel and distributed technology: systems and applications, 3(3):70–??, Fall 1995. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1995/pdf/h30070.pdf.

Anonymous:1995:BRd

[Ano95h] Anonymous. Book reviews. IEEE parallel and distributed technology: systems and applications, 3(4):89–??, Winter 1995. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1995/pdf/h40089.pdf.

Anonymous:1995:CCPa

[Ano95i] Anonymous. Calendar/call for papers. IEEE parallel and distributed technology: systems and applications, 3(1):94–95, Spring 1995. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1995/pdf/h10094.pdf.

Anonymous:1995:CCPb

[Ano95j] Anonymous. Calendar/call for papers. IEEE parallel and distributed technology: systems and applications, 3(4):100–??, Winter 1995. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1995/pdf/h40100.pdf.

Anonymous:1995:CSMa

[Ano95k] Anonymous. Computer Society membership info. IEEE parallel and distributed technology: systems and applications, 3(2):(inside back cover), Summer 1995. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Anonymous:1995:CSMb

[Ano95l] Anonymous. Computer Society membership info. IEEE parallel and distributed technology: systems and applications, 3(3):48, Fall 1995. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Anonymous:1995:CSMc

[Ano95m] Anonymous. Computer Society membership info. IEEE parallel and distributed technology: systems and applications, 3(4): (inside back cover), Winter 1995. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Anonymous:1995:NB

[Ano95n] Anonymous. New books.

IEEE parallel and distributed technology: systems and applications, 3(4):96–??, Winter 1995. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1995/pdf/h40096.pdf.

Anonymous:1995:NPa

[Ano95o] Anonymous. New products.

IEEE parallel and distributed technology: systems and applications, 3(1):86–??, Spring 1995. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1995/pdf/h10086.pdf.

Anonymous:1995:NPb

[Ano95p] Anonymous. New products. IEEE parallel and distributed technology: systems and applications, 3(2):72–??, Summer 1995. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1995/pdf/h20072.pdf.

Anonymous:1995:NPc

[Ano95q] Anonymous. New products. IEEE parallel and distributed technology: systems and applications, 3(3):76–??, Fall 1995. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1995/pdf/h30076.pdf.

Anonymous:1995:NPd

[Ano95r] Anonymous. New products. IEEE parallel and distributed technology: systems and applications, 3(4):97–??, Winter 1995. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1995/pdf/h40097.pdf.

Anonymous:1995:Na

[Ano95s] Anonymous. News. IEEE parallel and distributed technology: systems and applications, 3(2):76–??, Summer 1995. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1995/pdf/h20076.pdf.

Anonymous:1995:Nc

[Ano95t] Anonymous. News. IEEE parallel and distributed technology: systems and applications, 3(4):85–??, Winter 1995. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1995/pdf/h40085.pdf.

Anonymous:1995:RSSa

[Ano95u] Anonymous. Reader service/ subscription cards. IEEE parallel and distributed technology: systems and applications, 3(1): 96A–86B, Spring 1995. CO-DEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Anonymous:1995:RSSb

[Ano95v] Anonymous. Reader service/ subscription cards. IEEE parallel and distributed technology: systems and applications, 3(2): 88A–88B, Summer 1995. CO-DEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Anonymous:1995:RSSc

[Ano95w] Anonymous. Reader service/ subscription cards. IEEE parallel and distributed technology: systems and applications, 3(4): 80A–80B, Winter 1995. CO-DEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Anonymous:1995:RSSd

[Ano95x] Anonymous. Reader service/ subscription cards. IEEE parallel and distributed technology: systems and applications, 3(4): 104A–104B, Winter 1995. CO-DEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Anonymous:1996:R

[Ano96a] Anonymous. 1995 referees. IEEE parallel and distributed technology: systems and applications, 4(1):100, Spring 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1996/pdf/p1100.pdf.

Anonymous:1996:I

[Ano96b] Anonymous. 1996 index. IEEE parallel and distributed technology: systems and applications, 4(4):74–??, Winter 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1996/pdf/p4074.pdf.

Anonymous:1996:APIa

[Ano96c] Anonymous. Advertiser/product index. IEEE parallel and distributed technology: systems and applications, 4(1):56, Spring 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Anonymous:1996:APIb

[Ano96d] Anonymous. Advertiser/product index. IEEE parallel and distributed technology: systems and applications, 4(2):96, Summer 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Anonymous:1996:APIc

[Ano96e] Anonymous. Advertiser/product index. IEEE parallel and distributed technology: systems and applications, 4(3):(inside back cover), Fall 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Anonymous:1996:ASa

[Ano96f] Anonymous. Article summaries. IEEE parallel and distributed technology: systems and applications, 4(1):3–5, Spring 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1996/pdf/p1003.pdf.

Anonymous:1996:ASb

[Ano96g] Anonymous. Article summaries. IEEE parallel and distributed technology: systems and applications, 4(2):3–5, Summer 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1996/pdf/p2003.pdf.

Anonymous:1996:ASc

[Ano96h] Anonymous. Article summaries. IEEE parallel and distributed technology: systems and applications, 4(3):4–5, Fall 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1996/pdf/p3004.pdf.

Anonymous:1996:ASd

[Ano96i] Anonymous. Article summaries. IEEE parallel and distributed technology: systems and applications, 4(4):11–??, Winter 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1996/pdf/p4011.pdf.

Anonymous:1996:BRa

[Ano96j] Anonymous. Book reviews. IEEE parallel and distributed technology: systems and applications, 4(2):83–??, Summer 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1996/pdf/p2083.pdf.

Anonymous:1996:BRb

[Ano96k] Anonymous. Book reviews. IEEE parallel and distributed technology: systems and applications, 4(3):70–??, Fall 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1996/pdf/p3070.pdf.

Anonymous:1996:BRc

[Ano96l] Anonymous. Book reviews.

IEEE parallel and distributed technology: systems and applications, 4(4):78–??, Winter 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1996/pdf/p4078.pdf.

Anonymous:1996:CCPa

[Ano96m] Anonymous. Calendar/call for papers. IEEE parallel and distributed technology: systems and applications, 4(1):101–104, Spring 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1996/pdf/p1101.pdf.

Anonymous:1996:CCPb

[Ano96n] Anonymous. Calendar/call for papers. IEEE parallel and distributed technology: systems and applications, 4(2):94–96, Summer 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1996/pdf/p2094.pdf.

Anonymous:1996:CCPc

[Ano96o] Anonymous. Calendar/call for papers. IEEE parallel and distributed technology: systems and applications, 4(3):84–88, Fall 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1996/pdf/p3084.pdf.

Anonymous:1996:CCPd

[Ano96p] Anonymous. Calendar/call for papers. IEEE parallel and distributed technology: systems and applications, 4(4):87–??, Winter 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd1996/pdf/p4087.pdf.

Anonymous:1996:CSMa

[Ano96q] Anonymous. Computer Society membership info. IEEE parallel and distributed technology: systems and applications, 4(1): (inside back cover), Spring 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Anonymous:1996:CSMb

[Ano96r] Anonymous. Computer Society membership info. IEEE parallel and distributed technology: systems and applications, 4(2):(inside back cover), Summer 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Anonymous:1996:CSMc

[Ano96s] Anonymous. Computer Society membership info. IEEE parallel and distributed technology: systems and applications, 4(3):6, Fall 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Anonymous:1996:CWW

[Ano96t] Anonymous. Conferences: What will it take to sell a massive number of massively parallel machines? an IPPS ’96 panel report. IEEE parallel and distributed technology: systems and applications, 4(3):63– ??, Fall 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1996/pdf/p3063.pdf.

Anonymous:1996:ECS

[Ano96u] Anonymous. From the Editor-in-Chief: Strategic directions. IEEE parallel and distributed technology: systems and applications, 4(4):??, Winter 1996. CO-DEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Anonymous:1996:ISP

[Ano96v] Anonymous. Industry spotlight:

Personal privacy: an endangered species in the Information Age? IEEE parallel and distributed technology: systems and applications, 4(4):??, Winter 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Anonymous:1996:NBa

[Ano96w] Anonymous. New books. IEEE parallel and distributed technology: systems and applications, 4(1):92–??, Spring 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1996/pdf/p1092.pdf.

Anonymous:1996:NBb

[Ano96x] Anonymous. New books. IEEE parallel and distributed technology: systems and applications, 4(3):77–??, Fall 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd1996/pdf/p3077.pdf.

Anonymous:1996:NPa

[Ano96y] Anonymous. New products. IEEE parallel and distributed technology: systems and applications, 4(1):93–??, Spring 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd1996/pdf/p1093.pdf.

Anonymous:1996:NPb

[Ano96z] Anonymous. New products. IEEE parallel and distributed technology: systems and applications, 4(2):82–??, Summer 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd1996/pdf/p2082.pdf.

Anonymous:1996:NPc

[Ano96-27] Anonymous. New products. IEEE parallel and distributed technology: systems and applications, 4(3):78–??, Fall 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd1996/pdf/p3078.pdf.

Anonymous:1996:NPd

[Ano96-28] Anonymous. New products.

IEEE parallel and distributed technology: systems and applications, 4(4):85–??, Winter 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1996/pdf/p4085.pdf.

Anonymous:1996:Na

[Ano96-29] Anonymous. News. IEEE parallel and distributed technology: systems and applications, 4(1):98–??, Spring 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1996/pdf/p1098.pdf.

Anonymous:1996:Nb

[Ano96-30] Anonymous. News. IEEE parallel and distributed technology: systems and applications, 4(2):80–??, Summer 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1996/pdf/p2080.pdf.

Anonymous:1996:Nc

[Ano96-31] Anonymous. News. IEEE parallel and distributed technology: systems and applications, 4(3):81– ??, Fall 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1996/pdf/p3081.pdf.

Anonymous:1996:RSSa

[Ano96-32] Anonymous. Reader service/ subscription cards. IEEE parallel and distributed technology: systems and applications, 4(1): 104A–104A, Spring 1996. CO-DEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Anonymous:1996:RSSb

[Ano96-33] Anonymous. Reader service/ subscription cards — inside back cover. IEEE parallel and distributed technology: systems and applications, 4(2):96A–96A, Summer 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Anonymous:1996:RSSc

[Ano96-34] Anonymous. Reader service/ subscription cards — inside back cover. IEEE parallel and distributed technology: systems and applications, 4(3):88A–88A, Fall 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Bailey:1993:NPB

[BBDS93] D. H. Bailey, E. Barszcz, L. Dagum, and H. D. Simon. NAS parallel benchmark results. IEEE parallel and distributed technology: systems and applications, 1 (1):43–51, February 1993. CO-DEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Bass:1994:ADD

[BBH^{+}94] Julian M. Bass, Adrian R. Browne, Mohamed S. Hajji, Derek G. Marriott, Peter R. Croll, and Peter J. Fleming. Automating the development of distributed control software. IEEE parallel and distributed technology: systems and applications, 2(4):9–19, Winter 1994. CO-DEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Basten:1995:SAR

[BBV95] Twan Basten, Roland Bol, and Marc Voorhoeve. Simulating and analyzing railway interlockings in ExSpect. IEEE parallel and distributed technology: systems and applications, 3(3):50– 62, Fall 1995. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1995/pdf/h30050.pdf;   http:
//www.computer.org/concurrency/
pd1995/p3050abs.htm.

Bartoli:1995:GDD

[BCDP95] Alberto Bartoli, Paolo Corsini, Gianluca Dini, and Cosimo Antonio Prete. Graphical design of distributed applications through reusable components. IEEE parallel and distributed technology: systems and applications, 3(1):37–50, Spring 1995. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1995/pdf/h10037.pdf;   http:
//www.computer.org/concurrency/
pd1995/p1037abs.htm.

Boari:1995:ARD

[BCSL95] Maurelio Boari, Antonio Corradi, Cesare Stefanelli, and Letizia Leonardi. Adaptive routing for dynamic applications in massively parallel architectures. IEEE parallel and distributed technology: systems and applications, 3(1):61–74, Spring 1995. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd1995/pdf/h10061.pdf;   http:
                                        //www.computer.org/concurrency/
                                        pd1995/p1061abs.htm.

Blume:1994:ADP

[BEH^{+}94] William Blume, Rudolf Eigenmann, Jay Hoeflinger, David Padua, Paul Petersen, Lawrence Rauchwerger, and Peng Tu. Automatic detection of parallelism: a grand challenge for highperformance computing. IEEE parallel and distributed technology: systems and applications, 2(3):37–47, Fall 1994. CO-DEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Bell:1994:VWT

[Bel94] G. Bell. Visions: Why there won’t be apps: the problem with MPPs. IEEE parallel and distributed technology: systems and applications, 2(3):5–6, Fall 1994. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Browne:1995:VPD

[BHD^{+}95] James C. Browne, Syed I. Hyder, Jack Dongarra, Keith Moore, and Peter Newton. Visual programming and debugging for parallel computing. IEEE parallel and distributed technology: systems and applications, 3(1):75–83, Spring 1995. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1995/pdf/h10075.pdf;   http:
//www.computer.org/concurrency/
pd1995/p1075abs.htm.

Bokhari:1996:MCE

[Bok96] Shahid H. Bokhari. Multiphase complete exchange on Paragon, SP2, and CS-2. IEEE parallel and distributed technology: systems and applications, 4(3):45– 59, Fall 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1996/pdf/p3045.pdf;   http:
//www.computer.org/concurrency/
pd1996/p3045abs.htm.

Burleigh:1993:RDC

[Bur93] Scott Burleigh. ROME: distributing C++ object systems. IEEE parallel and distributed technology: systems and applications, 1(2):21–32, May 1993. CO-DEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Chang:1995:EPC

[CDH^{+}95] Sheue-Ling Chang, David Hung-Chang Du, Jenwei Hsieh, Rose P.

Tsang, and Mengjou Lin. Enhanced PVM communications over a high-speed LAN. IEEE parallel and distributed technology: systems and applications, 3(3):20–32, Fall 1995. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1995/pdf/h30020.pdf;   http:
//www.computer.org/concurrency/
pd1995/p3020abs.htm.

Chiola:1993:DSP

[CF93] Giovanni Chiola and Alois Ferscha. Distributed simulation of Petri nets. IEEE parallel and distributed technology: systems and applications, 1(3):33–50, August 1993. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Choudhary:1995:GEI

[CFS95] Alok Choudhary, Ian Foster, and Rick Stevens. Guest Editors’ introduction: Multimedia applications and high-performance computing. IEEE parallel and distributed technology: systems and applications, 3(2):2–3, Summer 1995. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1995/pdf/h20002.pdf.

Cox:1994:DSA

[CH94] Michael Cox and Pat Hanrahan. A distributed snooping algorithm for pixel merging. IEEE parallel and distributed technology: systems and applications, 2 (2):30–36, Summer 1994. CO-DEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Clark:1995:VBI

[Cla95] David Clark. Viewpoint: Buy it by the slice: The Microsoft vision for distributed computing. IEEE parallel and distributed technology: systems and applications, 3(4):8–11, Winter 1995. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1995/pdf/h40008.pdf.

Clark:1996:VTS

[Cla96] David Clark. Viewpoint: Toy Story — A technological leap. IEEE parallel and distributed technology: systems and applications, 4(2):6–8, Summer 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1996/pdf/p2006.pdf.

Crutcher:1995:MNT

[CLFZ95] Laurence A. Crutcher, Aurel A.

Lazar, Steven K. Feiner, and Michelle X. Zhou. Managing networks through a virtual world. IEEE parallel and distributed technology: systems and applications, 3(2):4–13, Summer 1995. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1995/pdf/h20004.pdf;   http:

                                        //www.computer.org/concurrency/
                                        pd1995/p2004abs.htm.

Choi:1996:TCD

[CLY96] Lynn Choi, Hock-Beng Lim, and Pen-Chung Yew. Techniques for compiler-directed cache coherence. IEEE parallel and distributed technology: systems and applications, 4(4):23–34, Winter 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd1996/pdf/p4023.pdf;   http:
                                        //www.computer.org/concurrency/
                                        pd1996/p4023abs.htm.

Calzarossa:1995:MTW

[CMM^{+}95] Maria Calzarossa, Luisa Massari, Alessandro Merlo, Mario Pantano, and Daniele Tessera. Medea: a tool for workload characterization of parallel systems. IEEE parallel and distributed technology: systems and applications, 3(4):72–80, Winter 1995. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd1995/pdf/h40072.pdf;   http:
                                        //www.computer.org/concurrency/
                                        pd1995/p4072abs.htm.

Crockett:1994:PPR

[CO94] Thomas W. Crockett and Tobias Orloff. Parallel polygon rendering for message-passing architectures. IEEE parallel and distributed technology: systems and applications, 2(2):17–28, Summer 1994. CODEN IPDTEX.

ISSN 1063-6552 (print), 1558-1861 (electronic).

Cownie:1994:VWM

[Cow94] J. Cownie. Visions: Why MPPs?

(parallel processing). IEEE parallel and distributed technology: systems and applications, 2(3):7– 8, Fall 1994. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Crowl:1994:HMP

[Cro94] Lawrence A. Crowl. How to measure, present, and compare parallel performance. IEEE parallel and distributed technology: systems and applications, 2(1):9–25, Spring 1994. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Chapman:1994:EHA

[CZM94] Barbara Chapman, Hans Zima, and Piyush Mehrotra. Extending HPF for advanced data-parallel applications. IEEE parallel and distributed technology: systems and applications, 2(3):59– 70, Fall 1994. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Dalpee:1993:BRV

[DC93] Michael J. Dalpee and T. James Cannaliato. Beyond RPC: the Virtual Network. IEEE parallel and distributed technology: systems and applications, 1(4): 41–57, November 1993. CO-DEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Disz:1995:DFC

[DEH^{+}95] Terrence L. Disz, Remy Evard, Mark W. Henderson, William Nickless, Robert Olson, Michael E. Papka, and Rick Stevens. Designing the future of collaborative science: Argonne’s Futures Laboratory. IEEE parallel and distributed technology: systems and applications, 3(2):14–21, Summer 1995. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1995/pdf/h20014.pdf;   http:
//www.computer.org/concurrency/
pd1995/p2014abs.htm.

Dongarra:1993:LAL

[Don93] J. Dongarra. Linear algebra libraries for high-performance computers: a personal perspective. IEEE parallel and distributed technology: systems and applications, 1(1):17–24, February 1993. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Du:1993:IAP

[Du93] Weichang Du. An intensional approach to parallel programming. IEEE parallel and distributed technology: systems and applications, 1(3):22–32, August 1993. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

El-Rewini:1993:PDS

[ERLS93] Hesham El-Rewini, Ted Lewis, and Bruce Shriver. Parallel and distributed systems — from theory to practice. IEEE parallel and distributed technology: systems and applications, 1(3):7–11, August 1993. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Ewing:1994:DCW

[ESM^{+}94] Richard E. Ewing, Robert C.

Sharpley, Derek Mitchum, Patrick O’Leary, and James S. Sochacki. Distributed computation of wave propagation models using PVM. IEEE parallel and distributed technology: systems and applications, 2(1):26–31, Spring 1994. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Feitelson:1995:PSM

[FCBH95] Dror G. Feitelson, Peter F. Corbett, Sandra Johnson Baylor, and Yarsun Hsu. Parallel I/O subsystems in massively parallel supercomputers. IEEE parallel and distributed technology: systems and applications, 3(3):33– 47, Fall 1995. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1995/pdf/h30033.pdf;   http:
//www.computer.org/concurrency/
pd1995/p3033abs.htm.

Fiedler:1994:VPH

[Fie94]  Stuart L. Fiedler.  Visions:

Putting HPC to work: a call to action. IEEE parallel and distributed technology: systems and applications, 2(4):5–7, Winter 1994. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Foster:1994:TPH

[Fos94] Ian Foster. Task parallelism and high-performance languages. IEEE parallel and distributed technology: systems and applications, 2(3):27–36, Fall 1994. CO-DEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Friesen:1995:DPP

[FYC95] Jerrold A. Friesen, Christine L. Yang, and Raymond E. Cline, Jr. DAVE: a plug-and-play model for distributed multimedia application development. IEEE parallel and distributed technology: systems and applications, 3(2):22–28, Summer 1995. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd1995/pdf/h20022.pdf;   http:
                                        //www.computer.org/concurrency/
                                        pd1995/p2022abs.htm.

Gorton:1995:OBM

[GGJ95] Ian Gorton, Jonathan P. Gray, and Innes Jelly. Object-based modeling of parallel programs. IEEE parallel and distributed technology: systems and applications, 3(2):52–63, Summer 1995. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd1995/pdf/h20052.pdf;   http:

//www.computer.org/concurrency/
pd1995/p2052abs.htm.

Grama:1993:IMS

[GGK93] Ananth Y. Grama, Anshul Gupta, and Vipin Kumar. Isoefficiency: measuring the scalability of parallel algorithms and architectures. IEEE parallel and distributed technology: systems and applications, 1(3):12–21, August 1993. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Gross:1994:TPH

[GOS94] Thomas Gross, David R. O’Hallaron, and Jaspal Subhlok. Task parallelism in a High Performance Fortran framework. IEEE parallel and distributed technology: systems and applications, 2(3):16– 26, Fall 1994. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Groselj:1993:BMG

[Gro93] Bojan Groselj. Bounded and minimum global snapshots. IEEE parallel and distributed technology: systems and applications, 1 (4):72–83, November 1993. CO-DEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Grimshaw:1993:DOP

[GSN93] Andrew S. Grimshaw, W. Timothy Strayer, and P. Narayan. Dynamic, object-oriented parallel processing. IEEE parallel and distributed technology: systems and applications, 1(2):33– 47, May 1993. CODEN IPDTEX.

ISSN 1063-6552 (print), 1558-1861 (electronic).

Grujic:1996:SSH

[GTM96] Aleksandra Grujić, Milo Tomašević, and Veljko Milutinović. A simulation study of hardwareoriented DSM approaches. IEEE parallel and distributed technology: systems and applications, 4(1):74–83, Spring 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd1996/pdf/p1074.pdf;   http:
                                        //www.computer.org/concurrency/
                                        pd1996/p1074abs.htm.

Gustafson:1994:VTO

[Gus94] J. Gustafson. Visions: Teraflops and other false goals. IEEE parallel and distributed technology: systems and applications, 2(2):5–6, Summer 1994. CO-DEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Gerasoulis:1993:PMC

[GY93] Apostolos Gerasoulis and Tao Yang. Pyrros: Mapping and compiling programs on scalable parallel architectures. IEEE parallel and distributed technology: systems and applications, 1 (3):81–82, August 1993. CO-DEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Hatcher:1994:GEI

[Hat94] Philip J. Hatcher. Guest Editor’s introduction: The impact of High Performance Fortran. IEEE parallel and distributed technology: systems and applications, 2(3):13–14, Fall 1994. CO-DEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Hansen:1994:GEI

[HCW94] Chuck Hansen, Tom Crockett, and Scott Whitman. Guest Editor’s introduction: Parallel rendering. IEEE parallel and distributed technology: systems and applications, 2(2):7, Summer 1994. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Heddaya:1993:MSS

[Hed93] Abdelsalam Heddaya. Memera:

Software shared memory on workstations and the CM-5. IEEE parallel and distributed technology: systems and applications, 1 (3):83–84, August 1993. CO-DEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Hedberg:1996:ISC

[Hed96a] S. Hedberg. Industry spotlight:

Combating computer viruses: IBM’s new computer immune system. IEEE parallel and distributed technology: systems and applications, 4(2):9–11, Summer 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1996/pdf/p2009.pdf.

Hedberg:1996:PPE

[Hed96b] S. Hedberg. Personal privacy:

an endangered species in the information age? IEEE parallel and distributed technology: systems and applications, 4(4):4–7, Winter 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Hedberg:1996:IST

[Hed96c] S. R. Hedberg. Industry spotlight: The telecommunications network of the new millennium. IEEE parallel and distributed technology: systems and applications, 4(1):6–8, Spring 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1996/pdf/p1006.pdf.

Hillis:1993:WFP

[Hil93] W. D. Hillis. Wrestling the future from the past: the transition to parallel computing. IEEE parallel and distributed technology: systems and applications, 1(1):6–7, February 1993. CO-DEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Hinrichs:1995:SCB

[Hin95]  Susan Hinrichs.  Simplifying

connection-based communication. IEEE parallel and distributed technology: systems and applications, 3(1):25–36, Spring 1995. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1995/pdf/h10025.pdf;   http:
//www.computer.org/concurrency/
pd1995/p1025abs.htm.

Hurson:1996:CMD

[HKSL96] A. R. Hurson, Krishna M. Kavi, Behrooz Shirazi, and Ben Lee. Cache memories for dataflow systems. IEEE parallel and distributed technology: systems and applications, 4(4):50–64, Winter 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1996/pdf/p4050.pdf;   http:
//www.computer.org/concurrency/
pd1996/p4050abs.htm.

Hammer:1994:DDR

[HLV^{+}94] Dieter K. Hammer, Erik J. Luit, Onno S. Van Roosmalen, Peter D. V. Van der Stok, and Jack P. C. Verhoosel. Dedos: a distributed real-time environment. IEEE parallel and distributed technology: systems and applications, 2(4):32–47, Winter 1994. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Huang:1994:CIP

[HM94] Chengchang Huang and Philip K.

McKinley. Communication issues in parallel computing across ATM networks. IEEE parallel and distributed technology: systems and applications, 2(4):73–86, Winter 1994. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Heath:1995:PPV

[HMR95] Michael T. Heath, Allen D.

Malony, and Diane T. Rover.

Parallel performance visualization: From practice to theory. IEEE parallel and distributed technology: systems and applications, 3(4):44–60, Winter 1995. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1995/pdf/h40044.pdf;   http:
//www.computer.org/concurrency/
pd1995/p4044abs.htm.

Harden:1995:SSB

[HRE^{+}95] Jim C. Harden, Donna S. Reese, Marlene B. Evans, Sudarshan Kadambi, Gregory J. Henley, Chuck E. Hudnall, and Cedell A. Alexander. In search of a standards-based approach to hybrid performance monitoring. IEEE parallel and distributed technology: systems and applications, 3(4):61–71, Winter 1995. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1995/pdf/h40061.pdf;   http:
//www.computer.org/concurrency/
pd1995/p4061abs.htm.

Irvin:1994:MSP

[IM94] R. Bruce Irvin and Barton P. Miller. Multiapplication support in a parallel-program performance tool. IEEE parallel and distributed technology: systems and applications, 2(1):40–50, Spring 1994. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Jadav:1995:DIH

[JC95] Divyesh Jadav and Alok Choudhary. Designing and implementing high-performance mediaon-demand servers. IEEE parallel and distributed technology: systems and applications, 3(2):29–39, Summer 1995. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1995/pdf/h20029.pdf;   http:
//www.computer.org/concurrency/
pd1995/p2029abs.htm.

Kahaner:1996:GBP

[Kah96] D. K. Kahaner. Global broadcast: Parallel computing in India. IEEE parallel and distributed technology: systems and applications, 4(3):7–11, Fall 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1996/pdf/p3007.pdf.

Krings:1994:ISR

[KKD94] Axel W. Krings, Roger M. Kieckhafer, and Jitender S. Deogun. Inherently stable real-time priority list dispatchers. IEEE parallel and distributed technology: systems and applications, 2 (4):49–59, Winter 1994. CO-DEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Kremien:1993:SAL

[KKM93] Orly Kremien, Jeff Kramer, and Jeff Magee. Scalable, adaptive load sharing for distributed systems. IEEE parallel and distributed technology: systems and applications, 1(3):62–70, August 1993. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Kotz:1995:FSW

[KN95] David Kotz and Nils Nieuwejaar. File-system workload on a scientific multiprocessor. IEEE parallel and distributed technology: systems and applications, 3(1):51–60, Spring 1995. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1995/pdf/h10051.pdf;   http:
//www.computer.org/concurrency/
pd1995/p1051abs.htm.

Kaddoura:1995:PUC

[KOR95] Maher Kaddoura, Chao-Wei Ou, and Sanjay Ranka. Partitioning unstructured computational graphs for nonuniform and adaptive environments. IEEE parallel and distributed technology: systems and applications, 3(3):63– 69, Fall 1995. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1995/pdf/h30063.pdf;   http:
//www.computer.org/concurrency/
pd1995/p3063abs.htm.

Lewis:1993:PTP

[LER93] Ted Lewis and Hesham El-Resini. Parallax: a tool for parallel program scheduling. IEEE parallel and distributed technology: systems and applications, 1(2):62– 72, May 1993. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Levesque:1994:APR

[Lev94] John M. Levesque. Applied Parallel Research’s xHPF system. IEEE parallel and distributed technology: systems and applications, 2(3):71, Fall 1994. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Loh:1996:HNT

[LHWS96] Peter Kok Keong Loh, Wen Jing Hsu, Cai Wentong, and Nadarajah Sriskanthan. How network topology affects dynamic load balancing. IEEE parallel and distributed technology: systems and applications, 4(3):25– 35, Fall 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1996/pdf/p3025.pdf;   http:
//www.computer.org/concurrency/
pd1996/p3025abs.htm.

Lin:1996:MMC

[Lin96] Yi-Bing Lin. Mobility management for cellular telephony networks. IEEE parallel and distributed technology: systems and applications, 4(4):65–73, Winter 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1996/pdf/p4065.pdf;   http:

                                        //www.computer.org/concurrency/
                                        pd1996/p4065abs.htm.

Lumetta:1993:DOP

[LML^{+}93] Steve Lumetta, Liam Murphy, Xiaoye Li, David C. Culler, and Ismail S. Khalil. Decentralized optimal power pricing: the development of a parallel program. IEEE parallel and distributed technology: systems and applications, 1 (4):23–31, November 1993. CO-DEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Loveman:1993:HPF

[Lov93] D. B. Loveman. High performance Fortran. IEEE parallel and distributed technology: systems and applications, 1(1): 25–42, February 1993. CO-DEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Li:1994:DAT

[LW94] Jingke Li and Michael Wolfe. Defining, analyzing, and transforming program constructs. IEEE parallel and distributed technology: systems and applications, 2(1):32–39, Spring 1994. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Miguel:1996:APN

[MABG96] José Miguel, Agustín Arruabarrena, Ramón Beivide, and José Angel Gregorio. Assessing the performance of the new IBM SP2 Communication Subsystem. IEEE parallel and distributed technology: systems and applications, 4(4):12–22, Winter 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1996/pdf/p4012.pdf;   http:
//www.computer.org/concurrency/
pd1996/p4012abs.htm.

May:1993:PPE

[MB93] John May and Francine Berman.

Panorama: portable, extensible debugger for parallel programs. IEEE parallel and distributed technology: systems and applications, 1(3):82, August 1993. CO-DEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Mackerras:1994:EDC

[MC94] Paul Mackerras and Brian Corrie. Exploiting data coherence to improve parallel volume rendering. IEEE parallel and distributed technology: systems and applications, 2(2):8–16, Summer 1994. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Messina:1993:CSC

[Mes93] P. Messina. The concurrent supercomputing consortium: Year 1. IEEE parallel and distributed technology: systems and applications, 1(1):9–16, February 1993. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Mechoso:1994:ASS

[MFS94] Carlos R. Mechoso, John D.

Farrara, and Joseph A. Spahr.

Achieving superlinear speedup on a heterogeneous, distributed system. IEEE parallel and distributed technology: systems and applications, 2(2):57–61, Summer 1994. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Mraz:1994:RVP

[Mra94] Ronald Mraz. Reducing the variance of point-to-point transfers for parallel real-time programs. IEEE parallel and distributed technology: systems and applications, 2(4):20–31, Winter 1994. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Morton:1996:MDO

[MT96] Donald J. Morton and John M. Tyler. Minimizing development overhead with partial parallelization. IEEE parallel and distributed technology: systems and applications, 4(3):15–24, Fall 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1996/pdf/p3015.pdf;   http:
//www.computer.org/concurrency/
pd1996/p3015abs.htm.

Mehra:1995:SWG

[MW95] Pankaj Mehra and Benjamin Wah. Synthetic workload generation for load balancing experiments. IEEE parallel and distributed technology: systems and applications, 3(3):4–19, Fall 1995. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1995/pdf/h30004.pdf;   http:
//www.computer.org/concurrency/
pd1995/p3004abs.htm.

Morton:1995:LLP

[MWO95] Don Morton, Kefei Wang, and David O. Ogbe. Lessons learned in porting Fortran/PVM code to the Cray T3D. IEEE parallel and distributed technology: systems and applications, 3(1):4–11, Spring 1995. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1995/pdf/h10004.pdf;   http:
//www.computer.org/concurrency/
pd1995/p1004abs.htm.

Netzer:1993:AML

[NX93] Robert H. B. Netzer and Jian Xu.

Adaptive message logging for incremental program replay. IEEE parallel and distributed technology: systems and applications, 1 (4):32–39, November 1993. CO-DEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Pande:1994:CFP

[PAM94] Santosh S. Pande, Dharma P.

Agrawal, and Jon Mauney. Compiling functional parallelism on distributed-memory systems. IEEE parallel and distributed technology: systems and applications, 2(1):64–76, Spring 1994. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Pancake:1993:GEI

[Pan93] Cherri M. Pancake. Guest Editor’s introduction: The changing face of supercomputing. IEEE parallel and distributed technology: systems and applications, 1 (4):12–15, November 1993. CO-DEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Peterson:1994:BET

[PC94] Gregory D. Peterson and Roger D. Chamberlain. Beyond execution time: Expanding the use of performance models. IEEE parallel and distributed technology: systems and applications, 2 (2):37–49, Summer 1994. CO-DEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Pernice:1996:RPP

[Per96]  Michael Pernice.  Review of

“PVM: Parallel Virtual Machine. A User’s Guide and Tutorial for Networked Parallel Computing”. IEEE parallel and distributed technology: systems and applications, 4(1):84, Spring 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd1996/pdf/p1084.pdf.

Pfister:1996:VCC

[Pfi96] G. F. Pfister. Viewpoint: Clusters of computers for commercial processing: the invisible architecture. IEEE parallel and distributed technology: systems and applications, 4(3):12– 14, Fall 1996. CODEN IPDTEX.

ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1996/pdf/p3012.pdf.

Parashar:1996:CTP

[PH96] Manish Parashar and Salim Hariri. Compile-time performance prediction of HPF/Fortran 90D. IEEE parallel and distributed technology: systems and applications, 4(1):57–73, Spring 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1996/pdf/p1057.pdf;   http:
//www.computer.org/concurrency/
pd1996/p1057abs.htm.

Ponnusamy:1995:SID

[PHD^{+}95] Ravi Ponnusamy, Yuan-Shin Hwang, Raja Das, Joel H. Saltz, Alok Choudhary, and Geoffrey Fox. Supporting irregular distributions using data-parallel languages. IEEE parallel and distributed technology: systems and applications, 3(1):12–24, Spring 1995. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1995/pdf/h10012.pdf;   http:
//www.computer.org/concurrency/
pd1995/p1012abs.htm.

Plank:1994:ICC

[PL94] James S. Plank and Kai Li.

ickp: a consistent checkpointer for multicomputers. IEEE parallel and distributed technology: systems and applications, 2(2):

62–67, Summer 1994. CO-DEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Pavlakos:1993:VMS

[PSM93] Constantine J. Pavlakos, Larry A. Schoof, and John F. Mareda. A visualization model for supercomputing environments. IEEE parallel and distributed technology: systems and applications, 1 (4):16–22, November 1993. CO-DEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Pancake:1995:GEI

[PSY95] Cherri M. Pancake, Margaret L. Simmons, and Jerry C. Yan. Guest Editors’ introduction: Performance evaluation tools for parallel and distributed systems. IEEE parallel and distributed technology: systems and applications, 3(4):14–20, Winter 1995. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd1995/pdf/h40014.pdf.

Protic:1996:DSM

[PTM96] Jelica Protić, Milo Tomašević, and Veljko Milutinović. Distributed shared memory: Concepts and systems. IEEE parallel and distributed technology: systems and applications, 4(2):63–79, Summer 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd1996/pdf/p2063.pdf;   http:

//www.computer.org/concurrency/
pd1996/p2063abs.htm.

Quinn:1994:ECa

[Qui94a] Michael J. Quinn. From the Editor-in-Chief. IEEE parallel and distributed technology: systems and applications, 2(2):4, Summer 1994. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Quinn:1994:ECb

[Qui94b] Michael J. Quinn. From the Editor-in Chief. IEEE parallel and distributed technology: systems and applications, 2(3):2–4, Fall 1994. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Quinn:1994:ECL

[Qui94c] Michael J. Quinn. From the Editor-in-Chief: Looking back and looking ahead. IEEE parallel and distributed technology: systems and applications, 2(4):4, Winter 1994. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Quinn:1995:R

[Qui95a] Michael J. Quinn. 1994 referees. IEEE parallel and distributed technology: systems and applications, 3(2):87–88, Summer 1995. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Quinn:1995:CSV

[Qui95b] Michael J. Quinn. Conferences: Supercomputing ’94: The view from the exhibition floor. IEEE parallel and distributed technology: systems and applications, 3(1):84–85, Spring 1995. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1995/pdf/h10084.pdf.

Reed:1995:CPG

[RCC^{+}95] Dan Reed, Charles Catlett, Alok Choudhary, David Kotz, and Marc Snir. Conferences: Parallel I/O: Getting ready for prime time. IEEE parallel and distributed technology: systems and applications, 3(2):64–71, Summer 1995. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1995/pdf/h20064.pdf.

ReeseHedberg:1995:PSD

[Ree95] S. Reese Hedberg. Parallelism speeds data mining. IEEE parallel and distributed technology: systems and applications, 3(4):3–6, Winter 1995. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1995/pdf/h40003.pdf.

Rosenblum:1995:CCS

[RHWG95] Mendel Rosenblum, Stephen A. Herrod, Emmett Witchel, and Anoop Gupta. Complete computer system simulation: The SimOS approach. IEEE parallel and distributed technology: systems and applications, 3(4):34–43, Winter 1995. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1995/pdf/h40034.pdf;   http:
//www.computer.org/concurrency/
pd1995/p4034abs.htm.

Rover:1993:VVI

[Rov93] Diane T. Rover. Vista: Visualization and instrumentation of scalable multicomputer applications. IEEE parallel and distributed technology: systems and applications, 1(3):83, August 1993. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Radia:1993:PVN

[RP93] Sanjay Radia and Jan Pachl.

The per-process view of naming and remote execution. IEEE parallel and distributed technology: systems and applications, 1 (3):71–80, August 1993. CO-DEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Reeves:1994:GEI

[RS94] Douglas S. Reeves and Kang G.

Shin. Guest Editors’ introduction: Parallel and distributed real-time computing. IEEE parallel and distributed technology: systems and applications, 2(4):8, Winter 1994. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Sistare:1994:SDM

[SAB^{+}94] Steve Sistare, Don Allen, Rich Bowker, Karen Jourdenais, Josh Simons, and R. Title. A scalable debugger for massively parallel message-passing programs. IEEE parallel and distributed technology: systems and applications, 2 (2):50–56, Summer 1994. CO-DEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Schuster:1994:PPG

[Sch94] Vincent J. Schuster. PGHPF from the Portland Group. IEEE parallel and distributed technology: systems and applications, 2(3):72, Fall 1994. CO-DEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Sever:1994:DRC

[SD94] Manfred D. M. Sever and Gabriele M. T. D’Eleuterio. Distributed, real-time control of structurally flexible manipulators. IEEE parallel and distributed technology: systems and applications, 2(4):61–71, Winter 1994. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Sarukkai:1995:ASA

[SMB95] Sekhar R. Sarukkai, Pankaj Mehra, and Robert J. Block. Automated scalability analysis of message-passing parallel programs. IEEE parallel and distributed technology: systems and applications, 3(4):21–32, Winter 1995. CODEN IPDTEX.

ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1995/pdf/h40021.pdf;   http:
//www.computer.org/concurrency/
pd1995/p4021abs.htm.

Spoerl:1994:SHC

[Spo94] James R. Spoerl. SofTech’s HPF compiler. IEEE parallel and distributed technology: systems and applications, 2(3):75, Fall 1994. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Secosky:1994:CCL

[SSBK94] Jason Secosky, Allan Snavely, Francine Berman, and Tom Kitchens. Conferences: Cruising the LeMans FunRun. IEEE parallel and distributed technology: systems and applications, 2 (2):68–71, Summer 1994. CO-DEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Stone:1993:MRO

[SSHT93] Janice M. Stone, Harold S. Stone, Philip Heidelberger, and John Turek. Multiple reservations and the Oklahoma update. IEEE parallel and distributed technology: systems and applications, 1 (4):58–71, November 1993. CO-DEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Schaeffer:1993:EMD

[SSLP93] Jonathan Schaeffer, Duane Szafron, Greg Lobe, and Ian Parsons. The Enterprise model for developing distributed applications. IEEE parallel and distributed technology: systems and applications, 1(3):85–96, August 1993. CO-DEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Sahni:1996:PMK

[ST96] Sartaj Sahni and Venkat Thanvantri. Performance metrics: Keeping the focus on runtime. IEEE parallel and distributed technology: systems and applications, 4(1):43–56, Spring 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1996/pdf/p1043.pdf;   http:
//www.computer.org/concurrency/
pd1996/p1043abs.htm.

Sun:1996:PPC

[SZ96] Xian-He Sun and Jianping Zhu. Performance prediction: a case study using a scalable shared-virtual-memory machine. IEEE parallel and distributed technology: systems and applications, 4(4):36–49, Winter 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1996/pdf/p4036.pdf;   http:
//www.computer.org/concurrency/
pd1996/p4036abs.htm.

Taylor:1995:SWM

[Tay95a] D. Taylor. Focus: Stormy weather main suspect in dinosaur disappearance. IEEE parallel and distributed technology: systems and applications, 3(4):12–13, Winter 1995. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1995/pdf/h40012.pdf.

Taylor:1995:CCPa

[Tay95b] Dennis Taylor. Calendar/call for papers. IEEE parallel and distributed technology: systems and applications, 3(2):84–86, Summer 1995. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1995/pdf/h20084.pdf.

Taylor:1995:CCPb

[Tay95c] Dennis Taylor. Calendar/call for papers. IEEE parallel and distributed technology: systems and applications, 3(3):78–80, Fall 1995. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1995/pdf/h30078.pdf.

Taylor:1995:CCPc

[Tay95d] Dennis Taylor. Calendar/call for papers. IEEE parallel and distributed technology: systems and applications, 3(4):100–104, Winter 1995. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Taylor:1996:FAS

[Tay96a] D. Taylor.  Focus:  An-

cient sounds, modern technology. IEEE parallel and distributed technology: systems and applications, 4(2):12–14, Summer 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1996/pdf/p2012.pdf.

Taylor:1996:FDM

[Tay96b] D. Taylor. Focus: Distributed multimedia for a song. IEEE parallel and distributed technology: systems and applications, 4(4):8–9, Winter 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1996/pdf/p4008.pdf.

Taylor:1995:MVD

[TCKK95] Herb Taylor, Danny Chin, Stan Knight, and James Kaba. The magic video-on-demand server and real-time simulation system. IEEE parallel and distributed technology: systems and applications, 3(2):40–51, Summer 1995. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1995/pdf/h20040.pdf;   http:
//www.computer.org/concurrency/
pd1995/p2040abs.htm.

Tomlinson:1993:ESS

[TCMW93] Christine Tomlinson, Phil Cannata, Greg Meredith, and Darrel Woelk. The extensible services switch in Carnot. IEEE parallel and distributed technology: systems and applications, 1(2):16– 20, May 1993. CODEN IPDTEX.

ISSN 1063-6552 (print), 1558-1861 (electronic).

Tandiary:1996:BUI

[TKDA96] Fredy Tandiary, Suraj C. Kothari, Ashish Dixit, and E. Walter Anderson. Batrun: Utilizing idle workstations for large-scale computing. IEEE parallel and distributed technology: systems and applications, 4(2):41–48, Summer 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1996/pdf/p2041.pdf;   http:
//www.computer.org/concurrency/
pd1996/p2041abs.htm.

Tokoro:1993:AOD

[Tok93] Mario Tokoro. Apertos: An object-oriented, distributed, realtime operating system. IEEE parallel and distributed technology: systems and applications, 1(3):84, August 1993. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Taylor:1996:CGB

[TWRP96] Stephen Taylor, Jerrell R. Watts, Marc A. Rieffel, and Michael E. Palmer. The concurrent graph: Basic technology for irregular problems. IEEE parallel and distributed technology: systems and applications, 4(2):15–25, Summer 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1996/pdf/p2015.pdf;   http:

                                        //www.computer.org/concurrency/
                                        pd1996/p2015abs.htm.

Vanderlip:1994:PSV

[Van94] Jeff Vanderlip. Pacific Sierra’s VAST-HPF and VAST/77toHPF. IEEE parallel and distributed technology: systems and applications, 2(3):74, Fall 1994. CO-DEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Veen:1994:PHP

[Vee94] Arthur H. Veen. The Prepare HPF programming environment. IEEE parallel and distributed technology: systems and applications, 2(3):73, Fall 1994. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Wallach:1994:VTL

[Wal94] S. Wallach. Visions: Teraflops into laptops. IEEE parallel and distributed technology: systems and applications, 2(3):8–9, Fall 1994. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Walker:1996:FFM

[Wal96] D. W. Walker. Focus: Freemarket computing and the integration of computers into the global economic infrastructure. IEEE parallel and distributed technology: systems and applications, 4(3):60–62, Fall 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

                                        dlib.computer.org/pd/books/

pd1996/pdf/p3060.pdf.  Dis-

cusses the role of Java as a standard for the Web, and how payment for computing resources might be triggered by author bank account information stored in Java applets.

Wladawsky-Berger:1994:PAN

[WB94] I. Wladawsky-Berger. Visions:

Parallel applications: the next frontier for computer industry breakthroughs. IEEE parallel and distributed technology: systems and applications, 2(3):10– 11, Fall 1994. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Wilson:1996:UCP

[WB96] Gregory V. Wilson and Henri E.

Bal. Using the Cowichan problems to assess the usability of Orca. IEEE parallel and distributed technology: systems and applications, 4(3):36– 44, Fall 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1996/pdf/p3036.pdf;   http:
//www.computer.org/concurrency/
pd1996/p3036abs.htm.

Weiss:1993:OSM

[Wei93] Shlomo Weiss. Optimizing a superscalar machine to run vector code. IEEE parallel and distributed technology: systems and applications, 1(2):73–83, May 1993. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Wu:1996:TUT

[WFL96] C. E. Wu, H. Franke, and Yew-Huey Liu. Tools: a unified trace environment for IBM SP systems. IEEE parallel and distributed technology: systems and applications, 4(2):89–93, Summer 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

                                        dlib.computer.org/pd/books/
                                        pd1996/pdf/p2089.pdf.

Wilson:1993:GPC

[Wil93] G. V. Wilson. A glossary of parallel computing terminology. IEEE parallel and distributed technology: systems and applications, 1 (1):52–67, February 1993. CO-DEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Wholey:1994:TMP

[WSL94] Skef Wholey, Richard Shapiro, and David Loshin. Thinking Machines’ plans for HPF. IEEE parallel and distributed technology: systems and applications, 2(3):76, Fall 1994. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Xu:1996:MCO

[XH96] Zhiwei Xu and Kai Hwang. Modeling communication overhead: MPI and MPL performance on the IBM SP2. IEEE parallel and distributed technology: systems and applications, 4(1):9–24, Spring 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1996/pdf/p1009.pdf;   http:
//www.computer.org/concurrency/
pd1996/p1009abs.htm.

Yen:1993:ERS

[YLB93] I-Ling Yen, E. L. Leiss, and F. B. Bastani. Exploiting redundancy to speed up parallel systems. IEEE parallel and distributed technology: systems and applications, 1(3):51–60, August 1993. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Yonezawa:1993:ICO

[YMYT93] Akinori Yonezawa, Satoshi Matsuoka, Masahiro Yasugi, and Kenjiro Taura. Implementing concurrent object-oriented languages on multicomputers. IEEE parallel and distributed technology: systems and applications, 1(2):49–61, May 1993. CO-DEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Zhang:1994:SLS

[ZCC94] Xiaodong Zhang, Robert Castaneda, and Elisa W. Chan. Spin-lock synchronization on the Butterfly and KSR1. IEEE parallel and distributed technology: systems and applications, 2(1):51–63, Spring 1994. CO-DEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Zait:1996:BDP

[ZFV96] Mohamed Zaït, Daniela Florescu, and Patrick Valduriez. Benchmarking the DBS3 Parallel Query Optimizer. IEEE parallel and distributed technology: systems and applications, 4(2):26–40, Summer 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1996/pdf/p2026.pdf;   http:
//www.computer.org/concurrency/
pd1996/p2026abs.htm.

Zomaya:1996:PPR

[Zom96] Albert Y. Zomaya. Parallel processing for real time simulation: a case study. IEEE parallel and distributed technology: systems and applications, 4(2):49–62, Summer 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1996/pdf/p2049.pdf;   http:
//www.computer.org/concurrency/
pd1996/p2049abs.htm.

Zhang:1996:EDS

[ZYC96] Xiaodong Zhang, Yong Yan, and Robert Castañeda. Evaluating and designing software mutual exclusion algorithms on shared-memory multiprocessors. IEEE parallel and distributed technology: systems and applications, 4(1):25–42, Spring 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://

dlib.computer.org/pd/books/
pd1996/pdf/p1025.pdf;   http:
//www.computer.org/concurrency/
pd1996/p1025abs.htm.